OpenStudy (anonymous):

which fractions are equivalent to 25/45

OpenStudy (anonymous):

5/9 is one

OpenStudy (anonymous):

i need two

OpenStudy (anonymous):

do you want me to give you the awnsers

OpenStudy (anonymous):

50/90 one more and maybe can i have a metal

OpenStudy (anonymous):

5/9 and 50/90

OpenStudy (anonymous):

hey i just said that

OpenStudy (anonymous):

that isnt correct here are the awnsers 5/9 and 15/28 or 5/9 and 75/135

OpenStudy (anonymous):

50/90 is correct

OpenStudy (anonymous):

but that isnt an awnser

OpenStudy (anonymous):

ay it is i will show you

OpenStudy (anonymous):

i know but that is not an awnser under the awnser choices

OpenStudy (anonymous):

25/45 = 5/9 = 10/18 = 15/27

OpenStudy (anonymous):

thank you because this shawn guy wont tell me and he is being rude

OpenStudy (anonymous):

i have the awnser now dont need youto prove anything

OpenStudy (anonymous):

see ya

OpenStudy (anonymous):

woww.., you are smart!

OpenStudy (anonymous):

this is the one that is right 5/9 and 75/135

OpenStudy (anonymous):

who

OpenStudy (anonymous):

:( i was not trying to be rude

OpenStudy (anonymous):

it is ok sorry to

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

me to

OpenStudy (anonymous):

that was it the other awnser was wrong thanks

OpenStudy (anonymous):

i just was not under standing that you were saying not that it was wrong but that i needed to pick out of 2 thing sorry

OpenStudy (anonymous):

also may be can you help me on something

OpenStudy (anonymous):

it ok

OpenStudy (anonymous):

what

OpenStudy (anonymous):

What is the surface area of the composite figure? http://learn.flvs.net/webdav/assessment_images/educator_math2/v9/07_PP_18_153.jpg two hundred fifty-two square millimeters two hundred sixty-four square millimeters three hundred square millimeters three hundred twenty-four square millimeters

OpenStudy (anonymous):

can you plz help me with thething that i put up

OpenStudy (anonymous):

three hundred square millimeters

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

any other ones you need help on

OpenStudy (anonymous):

i need help with one more if ok

OpenStudy (anonymous):

Which describes independent events? Spinning two fours on a spinner divided into four numbered sections. Selecting two kings from a standard deck of cards by choosing a card at random, placing it in your pocket, then choosing the second card. Selecting two green marbles by choosing one from a bag at random, giving the first to a friend, then choosing another. Selecting the names of two siblings by choosing slips of paper from a hat at random, pinning the selected slip on a bulletin board, and then selecting another.

OpenStudy (anonymous):

5/9 = 10/18 = 15/27 = 20/36 = 25/45 = 30/54 = 35/63 = 40/72 = 45/81 = 50/90 = 55/99 = 60/108 = 65/117 = 70/126 = 75/135 = 80/144 = 85/153 = 90/162 = 95/171 = 100/180